# Learning to Synthesize Fault-Tolerant Quantum Circuits by Diagonalization

### v1

@@ Abstract
Compilation of quantum programs into circuits expressed with discrete gate sets is essential for fault-tolerant quantum computing. Optimal methods for discovering high-precision implementations of unitaries in discrete gate sets such as the Clifford+T gate set are intractable. Search-based synthesis methods, including Reinforcement Learning (RL) and simulated annealing, are promising as they empirically discover efficient implementations of low-depth unitaries. We leverage search-based methods to reduce the general unitary synthesis problem to one of synthesizing diagonal unitaries; a problem solvable efficiently in general and optimally in the single-qubit case. Relying on several improvements in deep learning architectures essential for quantum datasets, our RL-based approach demonstrates how search-based methods and mathematical decompositions can be combined to efficiently find high precision implementations of unitaries taken from an array of real quantum algorithms. On these benchmarks we observe up to an average of $94\%$ fewer non-Clifford gates compared to the Quantum Shannon Decomposition.
Index Terms:
@@
quantum compilation, reinforcement learning, unitary synthesis, interpretability
I Introduction
Recent small scale demonstrations of logical qubits [1, 2] bode well for the future of Fault Tolerant (FT) quantum computing. In this regime, circuits are expressed in a discrete gate set, usually a small set of Clifford gates (e.g. H, S, CZ) augmented with a non-Clifford gate (e.g. T or V) for universality. The Clifford+T gate set can approximate any unitary transformation [3] and it has been used as the target of many FT quantum compilation algorithms [4, 5, 6, 7, 8]. As executing non-Clifford gates in many FT frameworks requires expensive pre-processing in the form of magic state distillation [9, 10, 11], a common goal of code generation tools is minimizing the number of non-Clifford gates in the circuit. Besides circuit quality (i.e. gate count), the scalability of FT compilers with respect to circuit size and required precision are crucial.
@@ I Introduction
It is known that if each entry of a unitary are elements of the ring $\mathbb{Z}[i,\frac{1}{\sqrt{2}}]$, that unitary can be implemented exactly with the Clifford+T gate set [12, 13]. Unitaries which appear in real algorithms are rarely this well structured, so they must be approximately synthesized instead. Analytical decomposition [14, 15] methods for Clifford+T generation are universally applicable, but circuit quality may suffer at high precision. In restricted cases, optimal approximate implementations of unitaries in discrete gate sets can be efficiently found [16]. Beyond single-qubit rotations restricted to a single axis, finding optimal approximate implementations of unitaries in the Clifford+T gate set is intractable [17].
@@
Search-based circuit synthesis methods perform discrete searches through the space of possible quantum gate placements [18, 19, 20]. Since this problem can be expressed as a Markov Decision Process, Reinforcement learning (RL) is a natural fit [21, 22, 23, 24, 25]. Search based algorithms can discover more resource efficient circuits than analytical methods, but fail to converge when compiling high precision implementations of complex unitaries. This is because continuously parameterized single-qubit rotation operations typically correspond to long sequences of discrete gates (see Figure 1). To ensure high circuit quality, these algorithms should be used in practice along with analytical methods.
We provide two contributions which enable hybrid search- and analytical-based methods to compile complex unitaries:
  1. 1.
@@
    We bypass the need to transpile difficult rotations with search-based tools by loosening the synthesis objective from matrix inversion to diagonalization. This allows search-based methods to synthesize $n$-qubit unitaries up to $2^{n}$-$1$ continuous degrees of freedom. Analytical decompositions and optimal single-qubit techniques are used to implement complex rotations. Diagonalization produces high precision results where other search-based tools fail. This process is general; other search-based tools can be retrofitted for diagonalization.
  2. 2.
    We introduce modifications to standard neural network architectures that enable machine learning methods to better tackle problems involving quantum datasets. Our Deep Value Network for Unitaries (DVNU) approach advances the state-of-the-art for RL based synthesis methods by improving the complexity of circuits which can be synthesized using deep learning.
@@
Our data indicate that the relaxation to diagonalization along with our neural architectural changes enable efficient and high precision synthesis of unitaries from real quantum algorithms. Our approach uses RL as its fast inference time makes it practical for use in full stack quantum compilers.
Figure 1: Strategies for synthesizing unitaries into Clifford+T gates. Analytical decomposition approaches, such as the Quantum Shannon Decomposition are universally applicable but often produces poor results. Search-based methods are able to find very high quality (low gate count) approximations, but are intractable except at low precision. Diagonalization is a hybrid approach that enables high quality and high precision. gridsynth [16] is used to optimally decompose continuous $R_{Z}$ operations (colored boxes) into Clifford+T gates.
@@
The remainder of the paper is organized as follows: Section II discusses necessary background information relating to synthesis with discrete gate sets. Section III explains how matrix diagonalization can be framed in the language of synthesis. Section IV introduces neural network architectural developments natural for quantum datasets. Section V evaluates the new architecture on synthesis tasks and shows how our RL based approach can effectively synthesize many unitaries taken from partitioned quantum algorithms. Section VI discusses methods of scaling, interprets the learned value function, and argues for the utility of RL in unitary synthesis. Finally, section VII offers concluding remarks.
II Background
II-A Discrete and Fault-Tolerant Gate Sets
@@
Universal quantum computers must represent complex unitaries on $n$ qubits with high fidelity. Modern noisy machines offer universal basis gate sets, which compilers use to express unitary operators. The fidelity with which these machines express algorithms is highly dependent on the choice of gate set [26]. Universal gate sets typically consists of an entangling two-qubit gate (e.g. CNOT or CZ gates) along with a set of fixed and continuously parameterized single-qubit gates (e.g. $X$ and $R_{Z}(\theta)$).
@@
Quantum Error Correcting Codes encode the state of one or more logical qubits into the states of many physical qubits [27]. Some logical operations require only that the same physical operation be applied to each physical qubit. These transversal gates form closed finite groups and are not universal [28]. The Clifford group coincides with the transversal gate set of many codes of interest. For universal quantum computation, the logical gate set must contain a non-transversal operation. The T gate is often considered as, being a member of the third level of the Clifford Hierarchy, it can be executed through magic state injection and Clifford gate corrections [29]. This work focuses on approximately compiling unitaries into the Clifford+T gate set.
Continuous rotations such as the $R_{Z}(\theta)$ gate are often unavoidable in quantum algorithms. For some precision $\epsilon$, translating these operations into Clifford+T gates requires $O(\log\frac{1}{\epsilon})$ gates. Except in select cases, this is true regardless of the angle of rotation. $R_{Z}$ rotations can be approximated optimally to arbitrary precision using the gridsynth algorithm [16]. Other methods also exist to handle these rotations [30, 31, 32], but for the scope of this paper we consider only the optimal ancilla free case.
II-B Circuit Synthesis
@@
Say that we want to synthesize the unitary labeled “Target” in Figure 1. Any unitary can be analytically decomposed into a sequence of CNOT and general single qubit unitary gates. The Quantum Shannon Decomposition and its variants are able to produce circuit implementations which approach the asymptotic lower bound of $O(4^{n})$ CNOT and $R_{Z}$ gates [14, 15]. Although these analytical approaches can always be applied and allow for arbitrarily high precision, they often produce circuits which contain many non-Clifford gates (i.e. $R_{Z}$ rotations). These $R_{Z}$ rotations can be implemented in several ways, including optimal decomposition into Clifford+T gates using gridsynth.
@@
Other algorithms often adopt a search-based approach. Here, a guiding function is used to iteratively modify a circuit, often gate-by-gate, until the target unitary is implemented. Examples of such algorithms include Reinforcement Learning approaches [21, 22, 23, 24, 25]. simulated annealing [20], and several optimal and heuristic synthesis algorithms [18, 17, 19]. Empirically, these state-of-the-art methods are able to find very efficient implementations of unitaries. Search-based tools are restricted in that they only operate with discrete gate sets, and therefore have no direct way of handling continuous rotations such as $R_{Z}$ gates. For unitaries which contain these rotations, search-based tools are only able to find low precision implementations.
Analytical and search-based approaches both seek to directly invert the adjoint of the target unitary. We propose an alternative approach: diagonalization (Figure 1). The diagonalization process consists of a two-headed search which halts when the target’s adjoint is diagonalized (i.e. not fully inverted). Diagonalizing ensures that up to $12^{n}-$ 1 continuous $R_{Z}$ operations can be handled. Although not optimal, this approach is able to find high precision implementations of complex unitaries (unlike search-based inversion methods), and produces significantly higher quality outputs than pure analytical methods (see Table II). In the worst case, diagonalization acts exactly like the underlying search-based algorithm which implements it.
II-C Formalizing Search-Based Quantum Circuit Synthesis
The goal of a quantum circuit synthesis algorithm is to find a quantum circuit implementation of a target unitary matrix. This problem is typically framed as finding a circuit which inverts the adjoint of the target.
@@ II-C Formalizing Search-Based Quantum Circuit Synthesis
When solving the synthesis problem directly by matrix inversion, we assume that every circuit can be described as a sequence of primitive gates taken from a finite set $A\subset\mathbb{U}(2^{n})$. Every primitive gate has a known associated unitary $a\in A$. A circuit consisting of $t$ gates can therefore be associated with a unitary matrix
$C_{t}=a_{t}\times\dots\times a_{1}.$ (1)
@@
$d_{HS}(C_{t},U_{tar})=\sqrt{1-\frac{1}{4^{n}}|Tr(C_{t}U_{tar}^{\dagger})|^{2}}\leq\epsilon$ (2)
is satisfied. Here $\epsilon\ll 1$ is a hyperparameter which determines how precisely a target must be implemented. The distance condition in Equation 2 is invariant to global phase. This is a desirable property as quantum processes that differ only by a global phase cannot be distinguished by projective measurements [29].
II-D Synthesis as a Markov Decision Process
Figure 2: The agent observes the state of the environment $s_{t}$, uses the policy to commit to an action $a_{t}$, then receives a reward $r_{t}$. Model-based synthesis approaches can evaluate the value of many actions before committing to them.
Markov Decision Processes (MDPs) provide a framework for describing stochastic decision problems. Posing quantum circuit synthesis as an MDP has been demonstrated numerous times in the literature [21, 22, 23, 24, 25] We define the MDP as a tuple $(S,S_{I},S_{T},A,r,p)$. Here $S$ is a set of states, $S_{I}$ a set of initial states, $S_{T}$ a set of terminal states, $A$ a set of actions or gates, $r:S\times A\rightarrow\mathbb{R}$ a reward function, and $p$ a distribution which assigns a probability to next states given a state-action pair.

### v2

# High Precision Fault-Tolerant Quantum Circuit Synthesis by Diagonalization using Reinforcement Learning

@@ Abstract.
Resource efficient and high precision compilation of programs into quantum circuits expressed in Fault-Tolerant gate sets, such as the Clifford+T gate set, is vital for the success of quantum computing. Optimal analytical compilation methods are known for restricted classes of unitaries, otherwise the problem is intractable. Empirical search-based synthesis methods, including Reinforcement Learning and simulated annealing, can generate good implementations for a more extensive set of unitaries, but require trade-offs in approximation precision and resource use. We leverage search-based methods to reduce the general unitary synthesis problem to one of synthesizing diagonal unitaries; a problem solvable efficiently in general and optimally in the single-qubit case. We demonstrate how our approach improves the implementation precision attainable by Fault-Tolerant synthesis algorithms on an array of unitaries taken from real quantum algorithms. On these benchmarks, many of which cannot be handled by existing approaches, we observe an average of 95% fewer resource-intensive non-Clifford gates compared to the more general Quantum Shannon Decomposition. On a subset of algorithms of interest for future term applications, diagonalization can reduce T gate counts by up to 16.8% compared to other methods.
@@
Quantum Compilation, Reinforcement Learning, Unitary Synthesis, Fault Tolerance
ccs: Hardware Quantum error correction and fault toleranceccs: Computing methodologies Reinforcement learningccs: Software and its engineering Compilers
1. Introduction
Recent small scale demonstrations of logical qubits (Bluvstein et al., 2023; da Silva et al., 2024) signal significant progress towards the development of Fault-Tolerant (FT) quantum computers. In this regime, a quantum program must be translated, or transpiled, into a quantum circuit (sequence of instructions) which uses only operations from a FT gate set. These discrete sets of gates consist of Clifford gates (e.g. H, S, CNOT) augmented with at least one non-Clifford gate (e.g. T or V) for universality. The Clifford+T gate set is one such universal FT gate set and is the target of many FT quantum compilation algorithms (Green et al., 2013; Developers, 2024; Sivarajah et al., 2020; Kissinger and van de Wetering, 2020; Javadi-Abhari et al., 2024). As executing non-Clifford gates in many models of FT computation requires expensive pre-processing in the form of magic state distillation (Knill, 2004; Bravyi and Kitaev, 2005), a common goal of code generation tools is minimizing the number of non-Clifford gates used while transpiling some target program.
Figure 1. Tradeoffs for synthesis algorithms targeting the Clifford+T gate set. Approximation accuracy refers to the precision with which a target program is implemented. In the FT setting, this value must be high in order for program outputs to be meaningful. Resource efficiency refers to non-Clifford gate counts. As these gates are orders of magnitude more expensive than Clifford gates, synthesis algorithms should use as few non-Clifford gates as possible. Unitary domain refers to how general the class of unitaries that can be taken as input is. This ranges from exactly synthesizable unitaries and approximations of 1-qubit $R_{Z}$ rotations, to the most general class of all multi-qubit unitaries.
The function of any $n$-qubit quantum circuit is fully specified by a unitary matrix $U\in\mathbb{U}(2^{n})$. As illustrated in Figure 1, existing FT compilation algorithms must trade-off between the quality of the solution (e.g. T gate count or depth), implementation precision (i.e. approximation error) and generality (i.e. unitary domain). The approximation error of a circuit measures the closeness of an algorithm’s implementation to the intended target program. Although the specific level of approximation error required for high-fidelity representation depends on the algorithm, low-precision synthesis eliminates the ability of synthesis algorithms to act as transpilers for future-term algorithms (see Section 6).
@@ 1. Introduction
If every element of a unitary is in the ring $\mathbb{Z}[e^{i\pi/4},\frac{1}{2}]$, it can be implemented exactly with the Clifford+T gate set (Kliuchnikov et al., 2012; Giles and Selinger, 2013). Unitaries which appear in real algorithms are rarely this well structured, meaning approximation is necessary. Luckily, the universal Clifford+T gate set can approximate any unitary transformation (Kitaev, 1997). General analytical decomposition methods (Shende et al., 2006; De Vos and De Baerdemacker, 2016) paired with single-qubit Clifford+T synthesis algorithms can handle any unitary, but the resultant circuits often contain an exponential ($O(4^{N})$) number of gates. In restricted cases (i.e. single qubit $R_{Z}$ rotations), optimal approximate Clifford+T implementations of unitaries can be efficiently found (Ross and Selinger, 2016). Beyond single-qubit rotations restricted to a single axis, finding optimal approximate implementations of unitaries in the Clifford+T gate set is intractable (Gosset et al., 2013).
@@
Search-based circuit synthesis methods perform discrete searches through the space of possible quantum gate placements (Amy et al., 2013; Gheorghiu et al., 2022; Paradis et al., 2024). Search-based algorithms can discover more efficient circuits than analytical methods, but fail when low approximation error is needed. Because continuously parameterized unitary rotation operations typically correspond to very long sequences of discrete gates (see Figure 2), existing methods fail to converge when compiling high precision implementations of complex unitaries.
Thus, the state-of-the-art in compiling FT circuits can be summarized as follows: we can generate precise implementations of arbitrary unitaries using a large number of resources, and we can generate resource efficient but highly coarse approximations of unitaries. Our work advances search-based FT circuit compilation in both precision and generality:
  1. (1)
@@
    We improve the approximation precision of multi-qubit synthesis algorithms on unitaries taken from real benchmarks. The main idea is to diagonalize rather than invert unitaries during synthesis. This allows us to bypass the need to transpile difficult single-qubit rotations with search-based tools. Instead we use optimal analytical techniques to implement complex single-qubit rotations. Diagonalization produces high precision results when state-of-the-art search-based tools fail. This approach is general; other search-based tools can be retrofitted for diagonalization.
  2. (2)
    Our Diagonalizing Value Network for Unitaries (DVNU) approach advances the state-of-the-art for RL-based synthesis methods by broadening the scope of circuits which can be synthesized using deep learning. We introduce two novel ideas: i) a global-phase invariance operation which alleviates the requirement that networks themselves learn invariant representations; ii) unitary perturbations applied during learning which force neural networks to associate a learned value with a region instead of a point. These methods can be applied to any existing machine-learning-based workflow that targets datasets which consist of unitary matrices.
@@
Our data indicate that the relaxation to diagonalization along with our neural architectural changes enable efficient and high precision synthesis of unitaries from real quantum algorithms. Our RL-based approach leads to fast inference and it is practical for use as a component in a full stack quantum compiler. When compared with existing techniques, we improve precision by several orders of magnitude. Synthesis using diagonalization also reduces T gate counts by up to 16.8% compared to existing techniques.
@@
The remainder of the paper is organized as follows: Section 2 discusses necessary background information relating to synthesis with discrete gate sets. Section 3 explains how matrix diagonalization can be framed in the language of synthesis. Section 4 introduces neural network architectural developments natural for quantum datasets. Section 5 evaluates the new architecture on synthesis tasks and shows how our RL-based approach can effectively synthesize many unitaries taken from partitioned quantum algorithms. Section 6 discusses methods of scaling, interprets the learned value function, and argues for the utility of RL in unitary synthesis. Finally, Section 7 offers concluding remarks.
2. Background
@@
Universal quantum computers must implement complex unitaries (programs) with high fidelity. Noisy Intermediate Scale Quantum (NISQ) gate sets typically consist of an entangling two-qubit gate (e.g. CNOT or CZ gates) along with a set of fixed and continuously parameterized single-qubit gates (e.g. $X$ and $R_{Z}(\theta)$).
@@
FT quantum computing differs in that it relies on quantum error correction, which restricts the kinds of operations that can be used. Quantum error correcting codes encode the state of one or more logical qubits into many physical qubits (Shor, 1995). Some logical operations require only that the same physical operation be applied to each physical qubit. These transversal gates form closed finite groups and are not universal (Eastin and Knill, 2009). The Clifford group coincides with the transversal gate set of many codes of interest. For universal quantum computation, the logical gate set must contain a non-transversal operation. This additional non-transversal gate is typically assumed to be the T gate, as such this work will focus on the approximate compilation of unitaries into the Clifford+T gate set. However, this method is extensible to a number of other ”magic” gates.
The T gate can be executed in a fault-tolerant manner through magic state distillation and injection with Clifford gate corrections (Nielsen and Chuang, 2011). However, state distillation is extremely costly. Estimates show that up to $99\%$ of the resources on a quantum computer could be dedicated to the implementation of these operations (Fowler et al., 2012; O’Gorman and Campbell, 2017). Minimizing the number of T gates for a given implementation of a unitary is therefore an essential step in the realization of practical error corrected quantum computation.
2.1. Fault-Tolerant Unitary Synthesis
A quantum program can be described as a circuit, or sequence of quantum gates. A unitary (or quantum circuit) synthesis algorithm prepares a circuit which implements some target unitary matrix. This implementation can be exact (e.g. zero error or zero process distance) or approximate. Exact synthesis is only possible for a very restricted subset of programs (unitaries). Only if every element of a unitary is in the ring $\mathbb{Z}[e^{i\pi/4},\frac{1}{2}]$ can it be implemented exactly with the Clifford+T gate set (Kliuchnikov et al., 2012; Giles and Selinger, 2013). Unitaries which appear in real algorithms are rarely this well structured, meaning approximation is necessary.
Figure 2. Comparison of analytical, search-based, and diagonalization strategies for Clifford+T synthesis. Analytical decomposition approaches, such as the Quantum Shannon Decomposition, are universally applicable but often produce poor quality results. Search-based methods are able to find very high quality (low gate count) approximations, but are intractable except at low precision and low qubit count. Diagonalization is a hybrid approach that enables both high quality and high precision. gridsynth (Ross and Selinger, 2016) is used to optimally decompose continuous single-qubit $R_{Z}$ operations (colored boxes) into Clifford+T gates.
@@ 2.1. Fault-Tolerant Unitary Synthesis
Say that we want to synthesize the unitary labeled Target in Figure 2. In the most general solution, as a first step, the unitary is decomposed using non-FT gates so that the program is expressed as a sequence of CNOT and general single-qubit unitary gates (any $2\times 2$ unitary matrix). The Quantum Shannon Decomposition and its variants are able to produce circuit implementations which approach the asymptotic lower bound of $O(4^{n})$ CNOT and $R_{Z}$ gates (Shende et al., 2006; De Vos and De Baerdemacker, 2016). Other numerical-optimization-based synthesis approaches can generate far more efficient circuits at the expense of compilation time (Davis et al., 2020; Smith et al., 2023). Although these approaches can always be applied and allow for low approximation error, they still often produce circuits which contain many non-Clifford gates.
[…]
@@
Another option is to deploy a synthesis algorithm which uses a guiding function to iteratively modify a circuit, often gate-by-gate, until the target unitary is implemented. These search-based approaches generate more resource efficient circuits than the previously described generic analytical workflow. Examples of such algorithms include Reinforcement Learning approaches (Zhang et al., 2020; Moro et al., 2021; Chen et al., 2022; Alam et al., 2023; Rietsch et al., 2024), simulated annealing (Paradis et al., 2024), and several optimal and heuristic synthesis algorithms (Amy et al., 2013; Gosset et al., 2013; Gheorghiu et al., 2022). Empirically, these state-of-the-art methods are able to find very efficient implementations of multi-qubit unitaries so long as solutions require few (meaning on the order of tens of) gates. Search-based tools are restricted in that they only operate with discrete gate sets, thus they have no direct way of handling continuous rotations such as $R_{Z}$ gates. For unitaries which contain these rotations, which are ubiquitous in unitaries taken from real benchmarks of interest, search-based tools are only able to find low precision implementations for a subset of inputs.
2.2. Formalizing Unitary Synthesis
@@
Unitary synthesis is typically framed as finding a circuit which inverts the adjoint (conjugate transpose) of the target matrix. When solving the problem directly by matrix inversion, every $n$-qubit circuit can be described as a sequence of primitive gates taken from a finite set $A\subset\mathbb{U}(2^{n})$. Every primitive gate has a known associated unitary $a\in A$. Note that a gate’s unitary representation $a$ depends both on the gate’s function and the qubit it acts on (Nielsen and Chuang, 2011). A circuit consisting of $t$ gates can therefore be associated with a unitary matrix
(1) $C_{t}=a_{t}\times\dots\times a_{1}.$
@@
(2) $d_{HS}(C_{t},U_{tar})=\sqrt{1-\frac{1}{4^{n}}|Tr(C_{t}U_{tar}^{\dagger})|^{2}}\leq\epsilon$
is satisfied. Here $\epsilon\ll 1$ is a hyperparameter which determines how precisely a target must be implemented.
Synthesis methods that invert the adjoint of a target unitary using an FT gate set must find a (potentially very long) sequence of discrete gates which satisfies Equation 2. We propose an alternative approach: diagonalization (Figure 2). The diagonalization process consists of a two-headed search which halts when the target’s adjoint is diagonalized (i.e. not fully inverted). Diagonalizing ensures that up to $12^{n}-$ 1 continuous $R_{Z}$ operations can be handled. Although there is no guarantee of optimality, this approach is able to find high precision implementations of complex unitaries (unlike search-based inversion methods), and produces significantly higher quality outputs than pure analytical methods (see Table 2). In the worst case, diagonalization acts exactly like the underlying search-based algorithm which implements it. When both diagonalization and search-based methods fail, analytical rule-based methods may be necessary.
Table 1 compares various unitary synthesis approaches which can be used to target the Clifford+T gate set. Among search-based methods, the simulated annealing tool Synthetiq empirically finds better implementations of unitaries than other methods (Paradis et al., 2024). Even so, Synthetiq fails to produce solutions for high precision implementations of complex unitaries as the space of circuits is too large to search. Synthesizing complex unitaries to high precision requires a mechanism for handling continuous rotations; this is possible in both analytical and diagonalization approaches.
Method Qubits Category Precision Unitary Domain
gridsynth (Ross and Selinger, 2016) 1 Analytical - Approx. $R_{Z}$
Policy Iter. (Alam et al., 2023) 1 Search (RL) $10^{-2}$ Approx. $\mathbb{U}(2)$
DQN (Chen et al., 2022) 1-2 Search (RL) $10^{-3}$ Approx. $\mathbb{U}(2)$-$\mathbb{U}(4)$
Synthetiq (Paradis et al., 2024) 1-4 Search (SA) $10^{-3}$ Approx. $\mathbb{U}(2)$-$\mathbb{U}(16)$
DVNU (ours) 1-3 Search (RL) $10^{-6}$ and below Approx. $\cup$ Diagonalizable $\mathbb{U}(2)$-$\mathbb{U}(8)$
QSD (Shende et al., 2006) 1+ Analytical - $U\in\mathbb{U}(2^{n})$
Table 1. Unitary synthesis approaches for the Clifford+T gate set. Among RL approaches, our Diagonalizing Value Network for Unitaries (DVNU) handles more complex inputs and with diagonalization can synthesize targets to very high precision. In cases where unitaries can be exactly diagonalized, DVNU can produce circuit implementations up to arbitrarily high precision. Precision here is measured using Hilbert-Schmidt distance (Eq 2). The Unitary Domain column indicates what kinds of unitaries can be handled. Synthetiq uses simulated annealing and empirically outperforms other pure search-based tools at low precision. gridsynth optimally decomposes 1-qubit $R_{Z}$ unitaries to any precision. The QSD is an analytical method which can be used along with gridsynth to synthesize Clifford+T circuits.
2.3. Synthesis as a Markov Decision Process
Figure 3. A model-based Reinforcement Learning agent. The agent observes the state of the environment $s_{t}$, uses the policy to commit to an action $a_{t}$, then receives a reward $r_{t}$. Model-based synthesis approaches can evaluate the value of many actions before committing to them.
Markov Decision Processes (MDPs) provide a framework for describing stochastic decision problems. Posing quantum circuit synthesis as an MDP has been demonstrated numerous times (Zhang et al., 2020; Moro et al., 2021; Chen et al., 2022; Alam et al., 2023; Rietsch et al., 2024). We define the MDP as a tuple $(S,S_{I},S_{T},A,r,p)$. Here $S$ is a set of states, $S_{I}$ a set of initial states, $S_{T}$ a set of terminal states, $A$ a set of actions or gates, $r:S\times A\rightarrow\mathbb{R}$ a reward function, and $p$ a distribution which assigns a probability to next states given a state-action pair. We use the term state (or sometimes unitary state) to describe an RL state $s_{t}$, not a quantum mechanical state vector or density operator.